# Josephson effect and self-trapping in helicoidal spin-orbit coupled Bose-Einstein condensates with optical lattices

### v3

@@ IV Effects of helicoidal gauge potentials and optical lattices on self-trapping
In the case of JT oscillation, we have seen that population imbalance (Z(t)) oscillates with time about $Z=0$. Understandably, the oscillation of $Z$ about zero imbalance depends on the values of the different parameters of the system. We notice that there exists critical values of the parameters for observation of JT oscillation beyond which the oscillation does not occur about $Z=0$ and thus an asymmetry in the maximum population imbalance arises between the two states. Due to the asymmetry, the average value of population imbalance over one period becomes non-zero. This causes the so-called quantum mechanical self-trapping (QMST)[48]. In order to derive a condition for the occurrence of self-trapping, we recast Eqs. (12) and (13) in terms of Hamiltonian equation of motion, $\dot{Z}=-\partial H/\partial\varphi$ and $\dot{\varphi}=\partial H/\partial Z$ with the Hamiltonian given by
@@
Variation of population imbalance with time for chosen values of different parameters satisfying the MQST condition is shown in Fig. 5. It shows that the population imbalance oscillates about a non-zero value and the amplitude of oscillation increases with the decrease of helicoidal gauge potential ($\beta$) (left panel). This implies that the self-trapping increases with the increase of halicoidal gauge potential. In order to quantify the amount of self-trapping, we calculate average value ($\langle Z(t)\rangle$) of population imbalance over a period for different values of $\beta$ and see that $\langle Z(t)\rangle$ increases nonlinearly as $\beta$ increases(right panel).
Figure 5: Left panel: Variation of population imbalance with $t$ for $\beta=3.2$ (solid blue) and $\beta=4$ (solid black). Right panel: It gives $\langle Z(t)\rangle$ versus $\beta$. In both the panels, we take $\Lambda=1.817$, $\alpha_{s}=4$, $x_{0}=0.25$, $G_{-}=2.05$, $k_{n}=1.8$ and $\Delta=0$.
@@
In order to find the effects of nonlinear optical lattices on the self-trapping, we plot $Z(t)$ for two different lattice strengths, namely, $G_{-}=2.05$ (solid blue)and $G_{-}=3.05$ (solid black) in Fig 7. We know that the nonlinear optical lattice influences the phase velocity of $Z(t)$ oscillation and thus plays an important role in changing the frequency of the oscillation [20]. The variation of frequency of oscillation ($\nu$) for different values of NOLs’ strengths is clearly reflected in the left panel of Fig. 8. More specifically, $\nu$ increases as the difference of lattice strengths ($G_{-}$) increases. The population imbalance in this case also oscillates about $Z(t)\neq 0$ and the amplitude of oscillation increases with the decrease of $G_{-}$. We calculate $\langle Z(t)\rangle$ for different values of $G_{-}$. The result displayed in Fig. 7 (right panel) shows that the self-trapping first increases and then attains a saturated value if $G_{-}$ further increases. We have checked that the self-trapping decreases if we take as $\Delta\neq 0$ for particular values of lattice and SOC parameters (Fig. 8).
Figure 7: Left panel: Variation of population imbalance with $t$ for $G_{-}=2.05$ (solid blue)and $G_{-}=3.05$ (solid black). Right panel: It gives $\langle Z(t)\rangle$ versus $G_{-}$. In both the panels we take $\beta=2$, $\alpha_{s}=2$, $x_{0}=0.25$, $k_{n}=1.8$, $\Lambda=0.681$ and $\Delta=0$.
Figure 8: Left panel: Variation of population imbalance with $t$ for $\Delta=0$ (solid blue) and $\Delta=1.2$ (solid black). Right panel: It gives $\langle Z(t)\rangle$ versus $\Delta$. In both the panels we take $\beta=5$, $\alpha_{s}=0.2$, $x_{0}=0.25$, $k_{n}=0.8$, $\Lambda=1.81$ and $G_{-}=2.05$.
@@
The variation of atomic density($\rho_{j}$=$\lvert\psi_{j}\lvert^{2}$) with time during quantum mechanical self-trapping is shown in Fig. 9. It clearly shows that the maximum density of one component corresponds to the minimum of others. However, the values of maxima or minima of atomic densities of two components are not equal and thus result MQST.
@@
We consider a helicoidal spin-orbit coupled Bose-Einstein condensate embedded in optical lattices and studying the time evolution of population imbalance between two pseudo-spin 1/2 states. More specifically, we consider two important cases for the system, namely, the effects of (i) helicoidal gauge potential and (ii) nonlinear optical lattices on population imbalance in a spin-orbit coupled Bose-Einstein condensate. It is seen that the population imbalance oscillates in time about a point with zero population imbalance. The amplitude of oscillation depends on the initial phase difference between the two states. In close analogy with the phase dependent current between two superconductor separated by a thin layer we refer the oscillation of atomic imbalance between spin-orbit coupled states as Josephson-type (JT) oscillation. We have seen that the frequency of oscillation increases as helicoidal gauge potential becomes weaker. In presence of nonlinear optical lattices, the frequency of Josephson-type oscillation for a fixed helicoidal gauge potential further increases.
We have examined that the oscillation of population imbalance does not occur about zero imbalance if the parameters of the system do not satisfy the so-called macroscopic quantum mechanical self-trapping condition. We quantify the self-trapping by calculating average value of population imbalance for different values of (i) helicoidal gauge potential, (ii) nonlinear lattice parameters and (iii) inter and itra-atomic interaction strengths. We find that, for a particular value of helicoidal gauge potential, self-trapping grows appreciably with the change of atomic interaction and lattice parameters. However, we observe saturation in the self-trapping for higher values of atomic interactions.

### v4

@@ IV Effects of helicoidal gauge potentials and optical lattices on self-trapping
In the case of JT oscillation, we have seen that population imbalance ($Z(t)$) oscillates with time about $Z=0$. Understandably, the oscillation of $Z$ about zero imbalance depends on the values of the different parameters of the system. We notice that there exists critical values of the parameters for observation of JT oscillation beyond which the oscillation does not occur about $Z=0$ and thus an asymmetry in the maximum population imbalance arises between the two states. Due to the asymmetry, the average value of population imbalance over one period becomes non-zero. This is termed as quantum mechanical self-trapping (QMST) [56]. In order to derive a condition for the occurrence of self-trapping, we recast Eqs. (11) and (12) in terms of Hamiltonian equations of motion, $\dot{Z}=-\partial H/\partial\varphi$ and $\dot{\varphi}=\partial H/\partial Z$ with the Hamiltonian given by
@@
Variation of population imbalance with time for chosen values of different parameters satisfying the MQST condition is shown in the left panel of Fig. 5. It shows that the population imbalance oscillates about a non-zero value and the amplitude of oscillation increases with the decrease of helicoidal gauge potential ($\beta$) (left panel). This implies that the self-trapping increases with the increase of halicoidal gauge potential. In order to quantify the amount of self-trapping, we calculate average value ($\langle Z(t)\rangle$) of population imbalance over a period for different values of $\beta$ and see that $\langle Z(t)\rangle$ increases nonlinearly as $\beta$ increases(right panel).
Figure 5: Left panel: Variation of population imbalance with $t$ for $\beta=3.2$ (solid blue) and $\beta=4$ (solid black). Right panel: It gives $\langle Z(t)\rangle$ versus $\beta$. In both the panels, we take $Z(0)=0.95$, $\Lambda=1.817$, $\alpha_{s}=4$, $x_{0}=0.25$, $G_{-}=2.05$, $k_{l}=5$, $k_{n}=1.8$ and $\Delta=0$.
@@
In order to find the effects of nonlinear optical lattices on the self-trapping, we plot $Z(t)$ for two different lattice strengths, namely, $G_{-}=2.05$ (solid blue) and $G_{-}=3.05$ (solid black) in Fig. 7 (left panel). We know that the nonlinear optical lattice influences the phase velocity of $Z(t)$ oscillation and thus plays an important role in changing the frequency of the oscillation [20]. The variation of population imbalance for different values of NOLs’ strengths shown in the left panel of Fig. 7 clearly indicates that $Z(t)$ oscillates with time and the amplitude of oscillation increases as the difference of lattice strengths ($G_{-}$) decreases. The population imbalance in this case also oscillates about $Z(t)\neq 0$ indicating self-trapping. We calculate $\langle Z(t)\rangle$ for different values of $G_{-}$. The result displayed in Fig. 7 (right panel) shows that the self-trapping first increases and then attains a saturated value if $G_{-}$ further increases.
In the above discussion we take zero Zeeman splitting. With a view to check the effects of Zeeman splitting ($\Delta$), we plot in Fig. 8, the variation of population imbalance with time for two different values of $\Delta$. It is seen that amplitude of oscillation increases while the frequency deceases for non-zero Zeeman spliting (left panel). The values of $\langle Z(t)\rangle$ decreases as $\Delta$ increases (right panel). This implies that the self-trapping decreases if we take as $\Delta\neq 0$ for particular choices of lattice and SOC parameters.
Figure 7: Left panel: Variation of population imbalance with $t$ for $G_{-}=2.05$ (solid blue)and $G_{-}=3.05$ (solid black). Right panel: It gives $\langle Z(t)\rangle$ versus $G_{-}$. In both the panels we take $Z(0)=0.95$, $\beta=2$, $\alpha_{s}=2$, $x_{0}=0.25$, $k_{l}=5$, $k_{n}=1.8$, $\Lambda=0.681$ and $\Delta=0$.
Figure 8: Left panel: Variation of population imbalance with $t$ for $\Delta=0$ (solid blue) and $\Delta=1.2$ (solid black). Right panel: It gives $\langle Z(t)\rangle$ versus $\Delta$. In both the panels we take $Z(0)=0.95$, $\beta=5$, $\alpha_{s}=0.2$, $x_{0}=0.25$, $k_{l}=5$, $k_{n}=0.8$, $\Lambda=1.81$ and $G_{-}=2.05$.
@@
The variation of atomic density ($\rho_{j}$=$\lvert\psi_{j}\lvert^{2}$) with time during quantum mechanical self-trapping is displayed in Fig. 9. It shows that the maximum and minimum densities of the second component are clearly distinguishable while they are not clearly distinguishable in the first component. This implies that most of the atoms are trapped in the first component and maximum or minimum of atomic density are not clearly visible. However, a few of atoms are involved in the exchange processes and thus causing density variation in the second component.
@@
We consider a helicoidal spin-orbit coupled Bose-Einstein condensate embedded in optical lattices and studying the time evolution of population imbalance between two pseudo-spin 1/2 states. More specifically, we consider two important ingredients for the system, namely, the effects of (i) helicoidal gauge potential and (ii) nonlinear optical lattices on population imbalance in a spin-orbit coupled Bose-Einstein condensate. It is seen that the population imbalance oscillates in time about a point with zero population imbalance. The amplitude of oscillation depends on the initial phase difference between the two states. In close analogy with the phase dependent current between two superconductor separated by a thin layer we refer the oscillation of atomic imbalance between spin-orbit coupled states as Josephson-type (JT) oscillation. We have seen that the frequency of oscillation increases as helicoidal gauge potential becomes weaker. In presence of nonlinear optical lattices, the frequency of Josephson-type oscillation for a fixed helicoidal gauge potential further increases.
We have examined that the oscillation of population imbalance does not occur about zero imbalance if the parameters of the system do not satisfy the so-called macroscopic quantum mechanical self-trapping condition. In order to quantify self-trapping, we calculating average value of population imbalance for different values of (i) helicoidal gauge potential, (ii) nonlinear lattice parameters and (iii) inter and intra-atomic interaction strengths. We find that, for a particular value of helicoidal gauge potential, self-trapping grows appreciably with the change of atomic interaction and lattice parameters. However, we observe saturation in the self-trapping for higher values of atomic interactions. Thus we conclude by noting that both the BECs with uniform and non-uniform SOCs can exhibit oscillation of atomic population imbalance and self-trapping. However, frequency of oscillations decreases as the system approaches towards the non-uniform SOC BECs. We also observe that self-trapping can be increased as non-uniformity in SOC increases.